### v3

# MARVEL: Unlocking the Multi-Modal Capability of Dense Retrieval via Visual Module Plugin

Abstract
This paper proposes Multi-modAl Retrieval model via Visual modulE pLugin (MARVEL), which learns an embedding space for queries and multi-modal documents to conduct retrieval. MARVEL encodes queries and multi-modal documents with a unified encoder model, which helps to alleviate the modality gap between images and texts. Specifically, we enable the image understanding ability of the well-trained dense retriever, T5-ANCE, by incorporating the visual module’s encoded image features as its inputs. To facilitate the multi-modal retrieval tasks, we build the ClueWeb22-MM dataset based on the ClueWeb22 dataset, which regards anchor texts as queries, and exacts the related text and image documents from anchor-linked web pages. Our experiments show that MARVEL significantly outperforms the state-of-the-art methods on the multi-modal retrieval dataset WebQA and ClueWeb22-MM. MARVEL provides an opportunity to broaden the advantages of text retrieval to the multi-model scenario. Besides, we also illustrate that the language model has the ability to extract image semantics and partly map the image features to the input word embedding space. All codes are available at https://github.com/OpenMatch/MARVEL.
MARVEL: Unlocking the Multi-Modal Capability of Dense Retrieval via Visual Module Plugin
[…]
1 Introduction
With the growth of multimedia information on the Internet, search engines tend to return multi-modal retrieval results to better satisfy the user information need Tautkute et al. (2019); Zhu et al. (2023). The media information provides more vivid retrieval results, such as texts, images, videos, and more, which improves users’ experiences and even changes their browsing behaviors.
Figure 1: Retrieval Pipeline with Our MARVEL Model. MARVEL incorporates the visual module plugin, aiming to unlock the multi-modal capabilities of well trained dense retrieval model.
Multi-modal retrieval Bain et al. (2021); Awad et al. (2021); Arni et al. (2008); Chang et al. (2022) aims to return fusion results of images and texts to answer user questions. The task can be modeled using a divide-and-conquer pipeline Chang et al. (2022); Liu et al. (2023b) or universal dense retrieval Liu et al. (2023b). UniVL-DR Liu et al. (2023b) encodes queries and multi-modal documents into a universal embedding space for multi-model retrieval. However, this work encodes image features and texts using different encoders from CLIP Radford et al. (2021) and the separated text and image encoding leads to a modality gap in representing multi-modal documents. It makes UniVL-DR design an additional image verbalization method to alleviate the modality gap and also limits the text retrieval models Karpukhin et al. (2020); Xiong et al. (2021a); Zhan et al. (2021); Li et al. (2021b); Yu et al. (2021) to excel their advantages in multi-modal scenarios.
In this paper, we propose Multi-modAl Retrieval model via Visual modulE pLugin (MARVEL). As shown in Figure 1, MARVEL is based on the text retriever T5-ANCE Yu et al. (2023), regards the visual module as a plugin and pretrains the visual module with image-caption contrastive training for adaption. By incorporating a visual module into well-trained text retriever T5-ANCE, MARVEL seizes the opportunity to extend the benefits of unimodal learning to the multi-modal retrieval task.
To facilitate the multi-modal retrieval task, we build a large-scale benchmark, ClueWeb22-MM, based on the web page dataset, ClueWeb22 Overwijk et al. (2022). Following previous work in text retrieval Zhang et al. (2020); Xie et al. (2023), we regard the anchor text as a query and assume that its linked web page is related to the query. Subsequently, we extract image and text documents from these anchor-linked web pages. After processing, the ClueWeb22-MM encompasses over 90k queries, maintaining a scale comparable to existing benchmark WebQA Chang et al. (2022). Previous work Xie et al. (2023) demonstrates that the high-quality training signals from anchor-document pairs contribute to developing a state-of-the-art dense retrieval model.
Our experiments show that MARVEL outperforms all baseline models, achieving improvements of over 2% and 7%, in the main metric MRR, on WebQA Chang et al. (2022) and ClueWeb22-MM, respectively. The evaluation results indicate the effectiveness of MARVEL comes from the visual module plugin architecture, the visual module pretraining method, and the text matching knowledge learned by T5-ANCE. Our further analyses illustrate that the image representations encoded by the visual module can be easily captured by only finetuning the language model parameters. The training strategies guide the language model to assign more appropriate attention weights to image and text features, preventing the visual module from overfitting to the training signals. These encoded image representations not only inhabit the input embedding space for semantics alignment but also function as a kind of prompt.
2 Related Work
Existing dense retrieval models Karpukhin et al. (2020); Xiong et al. (2021a); Ren et al. (2021); Xiong et al. (2021b); Gao and Callan (2022); Luan et al. (2021); Khattab and Zaharia (2020) usually focus on retrieving text documents and modeling the relevance between queries and documents. They usually employ pretrained language models to encode queries and text documents into an embedding space, followed by a KNN search to retrieve candidate documents.
Unlike the text retrieval task, the multi-modal retrieval task Chang et al. (2022); Hannan et al. (2020); Singh et al. (2021); Talmor et al. (2021) aims to provide users with multi-modal documents that satisfy their information needs. Earlier work primarily focuses on building a divide-and-conquer pipeline for multi-modal retrieval Chang et al. (2022); Liu et al. (2023b); Escalante et al. (2008); Grubinger et al. (2008). In these models, retrievers individually search candidates from the document collections of different modalities and then use a reranking model to fuse the retrieval results, such as vision-language models Zhang et al. (2021). However, this approach usually struggles to fuse the retrieval results across different modalities Liu et al. (2023b). UniVL-DR Liu et al. (2023b) builds a universal multi-modal dense retrieval model. It encodes queries and multi-modal documents as embeddings and conducts retrieval, modality routing, and result fusion within a unified embedding space.
Representing images is also the core of multi-modal retrieval, aiming to alleviate the modality gap between images and texts. Existing work usually focuses on representing the images using captions and image features Liu et al. (2023b) with different encoding methods. BERT-style pretrained visual-language models Chen et al. (2019); Lu et al. (2019); Tan and Bansal (2019); Su et al. (2020); Li et al. (2019, 2021a); Cho et al. (2021); Hu et al. (2020); Wang et al. (2022) provide an opportunity to model the captions and image features using the same model. However, these visual-language models typically aim to align the semantics between image features and captions instead of learning representations for image documents. Thus they show less effectiveness in learning an embedding space for multi-modal retrieval Liu et al. (2023b).
Another way to facilitate the image document representations is using the visual-language models that focus on representation learning, such as CLIP Radford et al. (2021). It encodes image features and texts using different encoders. However, these approaches often only provide shallow interactions between texts and visual features. Thus, existing models Liu et al. (2023b) pay more attention to alleviating the modality gap between texts and images by the image verbalization method, aiming to bridge the modality gap between images and texts in the raw text space.
Recent advancements in multi-modal large language models Brown et al. (2020); Touvron et al. (2023) have introduced a novel approach to modeling multi-modality features. This approach incorporates a visual encoder module into large language models through a transformation layer Li et al. (2023); Alayrac et al. (2022); Liu et al. (2023a). These models extract image features using the visual encoder module of CLIP and then optimize the prompt tokens and transformation layer to map the encoded image embeddings to the raw input space of large language models Merullo et al. (2023); Lester et al. (2021). Such a visual encoder plugin method presents a unified modeling approach for handling image and text features. It not only enables the visual comprehension ability of large language models but also preserves their effectiveness by freezing their parameters.
3 Multi-Modal Retrieval Model via Visual Module Plugin (MARVEL)
In this section, we first describe the multi-modal retrieval (Sec. 3.1) and then introduce the model architecture of MARVEL (Sec. 3.2).
(a) Visual Module Adaption Pretraining.
(b) Modality-Balanced Language Model Finetuning. We follow previous work Liu et al. (2023b) and sample one image document and one text document from corresponding negative document collections.
Figure 2: The Architecture of Multi-modAl Retrieval model via Visual modulE pLugin (MARVEL). We first pretrain the visual modules using the image-caption alignment task (Figure 2(a)) and then finetune the language model to conduct multi-modal retrieval (Figure 2(b)).
3.1 Preliminary of Multi-Modal Retrieval
Given a query $q$, the retrieval task requires the dense retrieval models to search relevant documents from the document collection $\mathcal{D}$ to meet the information needs of users.
Previous dense retrieval models Karpukhin et al. (2020); Xiong et al. (2021a); Gao and Callan (2021); Yu et al. (2023) usually focus on the text retrieval task, which aims to model the relevance between user query $q$ and text documents $\mathcal{D}=\{d_{\text{Text}}^{1},...,d_{\text{Text}}^{m}\}$. They encode both query and the $i$-th document $d_{\text{Text}}^{i}$ using language models, such as BERT Devlin et al. (2019), RoBERTa Liu et al. (2019) and T5 Raffel et al. (2020):
$\vec{q}=\text{TextEncoder}(q);\vec{d}_{\text{Text}}^{i}=\text{TextEncoder}(d_{\text{Text}}^{i}).$ (1)
Different from text retrieval Nguyen et al. (2016); Thakur et al. (2021), the multi-modal retrieval task Chang et al. (2022) aims to return a fusion result of documents from the collection $\mathcal{D}$, which are from different modalities. The document collection $\mathcal{D}$ not only contains texts $\mathcal{T}=\{d_{\text{Text}}^{1},...,d_{\text{Text}}^{m}\}$, but also includes images =$\mathcal{I}=\{d_{\text{Image}}^{1},...,d_{\text{Image}}^{n}\}$.
The multi-model retrieval task requires retrievers to conduct relevance modeling, cross-modal matching, and modality fusion Liu et al. (2023b). Previous work Liu et al. (2023b) maps text and image documents in an embedding space for retrieval, encodes texts and images using different encoders, and tries to bridge the modality gap using image verbalization methods. However, this limits the capability of dense retrieval models, hindering the expansion of text matching knowledge for learning representations for multi-modal documents.
3.2 Universal Multi-Modal Encoding
We show the model architecture in Figure 2. Different from previous work Liu et al. (2023b), we can universally encode query $q$ and multi-model documents $\mathcal{D}=\{d_{\text{Text}}^{1},...,d_{\text{Text}}^{m},d_{\text{Image}}^{1},...,d_{\text{Image}}^{n}\}$ using one encoder, T5-ANCE-CLIP:
$\begin{split}\vec{q}&=\text{T5-ANCE-CLIP}(q);\\ \vec{d}_{\text{Text}}^{i}&=\text{T5-ANCE-CLIP}(d_{\text{Text}}^{i});\\ \vec{d}_{\text{Image}}^{i}&=\text{T5-ANCE-CLIP}(d_{\text{Image}}^{i}(I),d_{\text{Image}}^{i}(C)),\end{split}$ (2)
where $d_{\text{Image}}^{i}(I)$ and $d_{\text{Image}}^{i}(C)$ are the image feature and caption of the $i$-th image document $d_{\text{Image}}^{i}$.
Then we calculate the relevance score $f(q,d^{i})$ between query $q$ and the $i$-th document $d^{i}$ using cosine similarity:
$f(q,d^{i})=\cos(\vec{q},\vec{d}^{i}).$ (3)
Following this, we conduct KNN search Johnson et al. (2019) to retrieve multi-modal document candidates for the given query $q$.
Subsequently, we first introduce the visual module plugin architecture of our MARVEL model (Sec. 3.2.1). Then we adapt the visual module to T5-ANCE by pretraining the visual understanding module (Sec. 3.2.2). Finally, we finetune the parameters of T5-ANCE to learn an embedding space for multi-modal retrieval (Sec. 3.2.3).
3.2.1 Dense Retrieval with Visual Plugin
MARVEL starts from the T5-ANCE model Yu et al. (2023), which is a dense retrieval model that is well-trained using MS MARCO dataset Nguyen et al. (2016). Then we enable T5-ANCE by incorporating the visual module from the vision-language model, CLIP Radford et al. (2021), and conduct the T5-ANCE-CLIP model. We can use a universal encoder, T5-ANCE-CLIP, to encode texts, image features and image documents.
Specifically, we encode the image feature $I$ using the visual encoder of CLIP Radford et al. (2021) and get its encoded visual representation $\vec{h}^{I}$:
$\vec{h}^{I}=\text{CLIP}(I),$ (4)
This representation is obtained from the grid features of the last layer of the visual encoder of CLIP, and $\vec{h}^{I}=\{\vec{h}^{I}_{1},...,\vec{h}^{I}_{49}\}$. Here 49 is the number of patches. Then we follow the previous visual-language model Merullo et al. (2023) and use a linear transformation layer to adapt the visual features $\vec{h}^{I}_{i}$ into the embedding space of the inputs of dense retrieval model:
$\vec{I}_{i}=\text{Linear}(\vec{h}^{I}_{i}).$ (5)
Finally, we can feed these encoded image features $\vec{I}=\{\vec{I}_{1},...,\vec{I}_{49}\}$ as the ahead input embeddings $\vec{X}$ for T5-ANCE:
$\vec{X}=\vec{e}(\text{<start>});\vec{I}_{1};...;\vec{I}_{49};\vec{e}(\text{<end>});\vec{e}_{1};...;\vec{e}_{k},$ (6)
[…]
Different from these visual-language models Alayrac et al. (2022); Li et al. (2023); Liu et al. (2023a); Tsimpoukelli et al. (2021), our MARVEL model aims to bring the advance of text retrieval-based pretraining to multi-modal retrieval tasks by using the visual model plugin to bridge the modality gap between images and texts.
3.2.2 Visual Module Adaption Pretraining
In MARVEL, we adapt the visual understanding module to T5-ANCE by only pretraining the parameters of the visual module (Eq. 4) and the projection layer (Eq. 5). We follow Radford et al. (2021) and leverage the image-caption contrastive training loss $L_{\text{VM}}$ to pretrain the visual understanding module. The training loss utilizes the alignment between image features $I$ and captions $C$:
$L_{\text{VM}}=L_{\text{IC}}+L_{\text{CI}},$ (7)
where $L_{\text{IC}}$ and $L_{\text{CI}}$ are the dual direction training losses to regard image and caption as queries and then map them with corresponding caption and image, respectively:
$L_{\text{IC}}=-\log\frac{e^{f(I,C^{+})/\tau}}{e^{f(I,C^{+})/\tau}+\sum_{C^{-}\in\mathcal{D}_{C}^{-}}{e^{f(I,C^{-})/\tau}}},$ (8)
$L_{\text{CI}}=-\log\frac{e^{f(C,I^{+})/\tau}}{e^{f(C,I^{+})/\tau}+\sum_{I^{-}\in\mathcal{D}_{I}^{-}}{e^{f(C,I^{-})/\tau}}},$ (9)
where $\tau$ is the temperature used to scale the similarity score. $\mathcal{D}_{C}^{-}$ and $\mathcal{D}_{I}^{-}$ contain negative captions and negative images respectively, which are sampled from in-batch negatives.
3.2.3 Modality-Balanced Language Model Finetuning
During finetuning, we can freeze the parameters of the visual module (Eq. 4) and optimize other parameters of MARVEL. To enable the MARVEL model to learn a universal embedding space for both queries and multi-modal documents, we follow previous work Liu et al. (2023b) and employ modality-balanced hard negative training to alleviate the modality discrimination of retrieval models:
$\begin{split}&L_{\text{LM}}=-\log\frac{e^{f(q,d^{+})/\tau}}{e^{f(q,d^{+})/\tau}+\sum_{d^{-}\in\mathcal{D}^{-}}{e^{f(q,d^{-})/\tau}}}\\ &\propto-\underbrace{f(q,d^{+})/\tau}_{L_{\text{Align}}}+\log(\sum_{d^{-}\in\mathcal{D}^{-}}(\underbrace{e^{f(q,d_{\,\text{Image}}^{-})/\tau}}_{L_{\text{Image}}}+\underbrace{e^{f(q,d_{\,\text{Text}}^{-})/\tau}}_{L_{\text{Text}}})),\end{split}$ (10)
where $\mathcal{D}^{-}$ contains the same number of negative documents of image and text. $L_{\text{Align}}$ teaches models to align the query with related documents. $L_{\text{Text}}$ and $L_{\text{Image}}$ guide retrievers to choose the modality and make the embedding space uniform Liu et al. (2023b); Wang and Isola (2020); Chen et al. (2020).
4 Experimental Methodology
This section describes datasets, evaluation metrics, baselines and implementation details.
Dataset Modality #Doc #Query
Train Dev Test
WebQA Image 389,750 16,400 2,554 2,511
Text 787,697 15,366 2,446 2,455
Multi-Modal 1,177,447 31,766 5,000 4,966
ClueWeb22-MM Image 368,710 35,873 5,041 5,030
Text 363,508 36,155 4,959 4,970
Multi-Modal 732,218 72,028 10,000 10,000
Table 1: Data Statistics.
Dataset. During pretraining, we collect the image-caption pairs from ClueWeb22 Overwijk et al. (2022) to train the visual understanding module. More details of pretraining data are shown in Appendix A.1. Then two multi-modal retrieval datasets, WebQA and ClueWeb22-MM, are used for finetuning and evaluation. The data statistics are shown in Table 1.
WebQA is a multi-hop, multi-modal, open-domain question answering benchmark (Chang et al., 2022). The dataset contains images and passage snippets that are crawled from the general Web and Wikipedia. We follow previous work Liu et al. (2023b) to keep the same experimental settings to preprocess the dataset. Besides, we build a new multi-modal retrieval dataset, ClueWeb22-MM, based on ClueWeb22 Overwijk et al. (2022), which provides 10 billion web pages with rich information. We only retain web pages in English and build the ClueWeb22-MM dataset. We establish query-document relations by pairing anchors with their corresponding document Xie et al. (2023); Zhang et al. (2020). And then we regard the anchor texts as queries and extract image documents and text documents from the linked documents. More details of building the ClueWeb22-MM dataset are shown in Appendix A.3.
Evaluation Metrics. We use NDCG@10, MRR@10 and Recall@100 as evaluation metrics. Following previous work  Liu et al. (2023b); Nguyen et al. (2016), we regard MRR@10 as our main evaluation. MRR and NDCG are computed using the official scripts [^1] https://github.com/microsoft/MSMARCO-Passage-Ranking/blob/master/ms_marco_eval.py. Statistic significances are tested by permutation test with P$<0.05$.
Baselines. In our experiments, we follow previous work Liu et al. (2023b) to conduct baseline models and divide them into three groups: single modality retrieval, divide-and-conquer, and universal dense retrieval models.
Single Modality Retrieval. In our experiments, we represent image documents using captions and use several text retrieval models as baselines. BM25 Robertson et al. (2009) is widely used in text retrieval work, which conducts exact matches between queries and documents. DPR Karpukhin et al. (2020) is trained using NQ dataset Kwiatkowski et al. (2019) and uses a dual-encoder to encode queries and documents as dense vectors for retrieval. We start from vanilla BERT Devlin et al. (2019) and DPR Karpukhin et al. (2020) checkpoints and train the encoders using in-batch negatives to conduct BERT-DPR and NQ-DPR models. NQ-ANCE is also compared, which continuously trains NQ-DPR using hard negatives Xiong et al. (2021a). Besides, T5-ANCE Yu et al. (2023) and Anchor-DR Xie et al. (2023) are dense retrieval models that are trained on MS MARCO and ClueWeb22, respectively.
Divide-and-Conquer. The divide-and-conquer models retrieve image documents and text documents individually and then fuse the retrieval results. Following previous work Liu et al. (2023b), we use single modality retrievers, VinVL-DPR, CLIP-DPR and BM25, and fuse the retrieval results according to their unimodal rank reciprocals.
Setting Model WebQA ClueWeb22-MM
MRR@10 NDCG@10 Rec@100 MRR@10 NDCG@10 Rec@100
Single Modality (Text Only) BM25 53.75 49.60 80.69 40.81 46.08 78.22
DPR (Zero-Shot) 22.72 20.06 45.43 20.59 23.24 44.93
CLIP-Text (Zero-Shot) 18.16 16.76 39.83 30.13 33.91 59.53
Anchor-DR (Zero-Shot) 39.96 37.09 71.32 42.92 48.50 76.52
T5-ANCE (Zero-Shot) 41.57 37.92 69.33 45.65 51.71 83.23
BERT-DPR 42.16 39.57 77.10 38.56 44.41 80.38
NQ-DPR 41.88 39.65 78.57 39.86 46.15 83.50
NQ-ANCE 45.54 42.05 69.31 45.89 51.83 81.21
Divide-Conquer VinVL-DPR 22.11 22.92 62.82 29.97 36.13 74.56
CLIP-DPR 37.35 37.56 85.53 39.54 47.16 87.25
BM25 & CLIP-DPR 42.27 41.58 87.50 41.58 48.67 83.50
UnivSearch CLIP (Zero-Shot) 10.59 8.69 20.21 16.28 18.52 40.36
VinVL-DPR 38.14 35.43 69.42 35.09 40.36 75.06
CLIP-DPR 48.83 46.32 86.43 42.59 49.24 87.07
UniVL-DR $\text{}^{\dagger\mathsection}$ $\text{}^{\dagger\mathsection}$ $\text{}^{\dagger\mathsection}$ $\text{}^{\dagger\mathsection}$ $\text{}^{\dagger\mathsection}$ $\text{}^{\dagger\mathsection}$
MARVEL-DPR $\text{}^{\dagger}$ $\text{}^{\dagger}$ $\text{}^{\dagger}$ $\text{}^{\dagger}$ $\text{}^{\dagger}$ $\text{}^{\dagger}$
MARVEL-ANCE $\text{}^{\dagger\ddagger\mathsection}$ $\text{}^{\dagger\ddagger\mathsection}$ $\text{}^{\dagger\ddagger\mathsection}$ $\text{}^{\dagger\ddagger\mathsection}$ $\text{}^{\dagger\ddagger\mathsection}$ $\text{}^{\dagger\ddagger\mathsection}$
[…]
Universal Dense Retrieval. CLIP-DPR and VinVL-DPR employ the visual language models CLIP Radford et al. (2021) and VinVL Zhang et al. (2021) as image and text encoders and then are trained with in-batch negatives. UniVL-DR Liu et al. (2023b) is our main baseline model, which further uses modality-balanced hard negative to train text and image encoders and also utilizes the image verbalization method to bridge the modality gap between images and texts.
Implementation Details. In our experiments, we use T5-ANCE Yu et al. (2023) as our backbone language model, which is well-trained on the MS MARCO dataset Nguyen et al. (2016). Then we implement our MARVEL model by utilizing CLIP as the visual understanding module to empower the image understanding capability of T5-ANCE. The visual encoder is initialized with the clip-vit-base-patch32 checkpoint from OpenAI [^2] https://github.com/openai/CLIP. For MARVEL, we truncate queries, text documents and image captions to 128 tokens and set the max number of visual tokens to 49.
During training, we use AdamW Loshchilov and Hutter (2019) optimizer and set maximum training epoch=20, batch size=64, learning rate=$5e-6$, and the temperature hyperparameter $\tau=0.01$. We follow UniVL-DR Liu et al. (2023b) and conduct MARVEL-ANCE by starting from in-batch negative finetuned MARVEL-DPR, and continuously training MARVEL-DPR with modality-balanced hard negatives. These hard negatives are randomly sampled from the top 100 retrieved negatives using MARVEL-DPR. All models are evaluated per 500 steps and the early stop step is set to 5.
Model Modality WebQA ClueWeb22-MM
MRR@10 NDCG@10 Rec@100 MRR@10 NDCG@10 Rec@100
MARVEL-ANCE Text $\text{}^{\ddagger}$ $\text{}^{\ddagger\mathsection}$ $\text{}^{\ddagger\mathsection}$ $\text{}^{\dagger\ddagger\mathsection}$ $\text{}^{\dagger\ddagger\mathsection}$ $\text{}^{\ddagger\mathsection}$
Image $\text{}^{\dagger}$ $\text{}^{\dagger\ddagger}$ $\text{}^{\dagger\ddagger\mathsection}$ $\text{}^{\dagger\ddagger\mathsection}$ $\text{}^{\dagger\ddagger\mathsection}$ $\text{}^{\dagger\ddagger}$
Multi $\text{}^{\ddagger}$ $\text{}^{\dagger\ddagger}$ $\text{}^{\ddagger\mathsection}$ $\text{}^{\ddagger\mathsection}$ $\text{}^{\ddagger\mathsection}$ $\text{}^{\ddagger\mathsection}$
w/o CLIP Pretraining Text $\text{}^{\ddagger}$ $\text{}^{\ddagger}$ $\text{}^{\ddagger\mathsection}$ $\text{}^{\mathsection}$ $\text{}^{\mathsection}$ $\text{}^{\ddagger\mathsection}$
Image 65.17 66.69 94.64 $\text{}^{\ddagger\mathsection}$ $\text{}^{\ddagger\mathsection}$ 96.22
Multi 64.66 $\text{}^{\ddagger}$ 92.24 $\text{}^{\ddagger\mathsection}$ $\text{}^{\ddagger\mathsection}$ $\text{}^{\ddagger\mathsection}$ $\text{}^{\ddagger}$
w/o MS MARCO Pretraining Text 63.37 56.93 88.54 $\text{}^{\mathsection}$ $\text{}^{\mathsection}$ 91.27
Image 65.73 66.91 94.66 76.26 80.11 96.08
Multi 64.21 61.63 91.43 $\text{}^{\mathsection}$ $\text{}^{\mathsection}$ 92.52
w/o Prompt Text 63.86 $\text{}^{\ddagger}$ $\text{}^{\ddagger}$ 69.99 73.82 91.65
Image $\text{}^{\dagger\ddagger}$ $\text{}^{\dagger\ddagger}$ 94.42 76.07 80.14 $\text{}^{\dagger\ddagger}$
Multi $\text{}^{\ddagger}$ $\text{}^{\ddagger}$ $\text{}^{\ddagger}$ 54.20 61.79 $\text{}^{\ddagger}$
[…]
5 Evaluation Result
In this section, we first evaluate the performance of MARVEL and then conduct ablation studies. Subsequently, we explore the effectiveness of different visual and language model fusion methods and analyze the role of image features in the MARVEL. Some case studies are shown in Appendix A.6.
5.1 Overall Performance
The multi-modal retrieval performance of MARVEL and baseline models is shown in Table 2.
Overall, MARVEL significantly outperforms baseline models on all datasets by achieving more than 2% improvements on both datasets, demonstrating its advantages in multi-modal retrieval tasks. Compared with text retrieval models, MARVEL improves their performance, showing that the image features are crucial in the multi-modal retrieval task. Furthermore, these universal multi-modal dense retrievers, UniVL-DR and MARVEL, outperform divide-and-conquer models by alleviating the modality fusion problem Liu et al. (2023b). Compared with our main baseline UniVL-DR, MARVEL encodes queries and multi-modal documents using a universal encoder. Experimental results show that MARVEL significantly improves the retrieval effectiveness of UniVL-DR on both datasets, demonstrating its effectiveness in bridging the modality gap between images and texts.
5.2 Ablation Study
As shown in Table 3, we conduct ablation studies to explore the role of different modules of MARVEL in multi-modal retrieval. More ablation studies are shown in Appendix A.5.
In the comparison between MARVEL and MARVEL (w/o CLIP Pretraining), pretraining the visual understanding module shows its effectiveness by improving the performance on single/multi-modal retrieval tasks. It shows that the image-caption alignment relations provide some opportunities to adapt the visual module to the language model via pretraining. Subsequently, MARVEL also outperforms MARVEL (w/o MS MARCO Pretraining), especially on the text retrieval task. It demonstrates that MARVEL can broaden the advantage of text relevance modeling to the multi-modal retrieval task. To unify the multi-modal encoding, MARVEL follows previous work Hannan et al. (2020) uses prompt tokens to indicate the start and end positions of encoded image features (Eq. 6), aiming to distinguish the image features from text token embeddings. These image prompt tokens bring light improvements, illustrating their roles in multi-modal document representation.
5.3 Retrieval Effectiveness of Different Visual-Language Fusion Methods
In this experiment, we show the retrieval effectiveness of MARVEL on the WebQA dataset by using different modality fusion and finetuning methods.
Method Modality MRR@10 Rec@100
Text 51.75 84.37
CLIP-Sum Image 60.61 94.84
Multi 48.83 86.43
Text 51.84 85.06
MARVEL-Sum Image 58.09 93.13
Multi 35.03 79.00
Text 48.71 81.78
MARVEL-Concat Image 37.20 81.14
Multi 25.19 62.77
Text 54.28 85.80
MARVEL Image 60.81 93.55
Multi 55.58 88.50
Table 4: Retrieval Performance of the Models using Different Visual-Language Fusion Methods. The Sum and Concat methods are similar to previous work Liu et al. (2023b), which only replace the image caption encoder with T5-ANCE. The CLIP-Sum model is the CLIP-DPR model from previous work Liu et al. (2023b). All models are trained with in-batch negatives. MRR@10 is used to evaluate the retrieval effectiveness of all models.
Modality Fusion. Three kinds of visual-language fusion methods are compared in our experiments, including Sum, Concat and Plugin. For Sum and Concat methods, we encode the captions and image features separately as embeddings, then sum or concatenate these embeddings, followed by joint training of T5-ANCE and CLIP models with in-batch negatives. We show the experimental results in Table 4. MARVEL’s visual module plugin method outperforms other fusion methods. This highlights the effectiveness of utilizing pretrained attention heads of language models for extracting image semantics and fostering deeper interactions between image and text inputs. Our plugin method proves instrumental in mitigating the modality gap between texts and images, enabling MARVEL to better represent image documents by jointly modeling image captions and features.
Different from Liu et al. (2023b), we use T5-ANCE and CLIP as the text and image encoders, respectively. These models have different architectures and are pretrained on text retrieval and image-caption matching tasks. The multi-modal retrieval performance of CLIP-Sum decreases when we encode the image caption with a stronger retrieval model T5-ANCE (MARVEL-Sum) instead of CLIP. It demonstrates that incorporating an additional visual module into a well-trained dense retrieval model is still challenging for multi-modal retrieval. Notably, MARVEL provides a promising way to enable the image understanding ability of dense retrieval models by using the visual module plugin modeling method.
(a) Attention Weights.
(b) Attention Entropy.
Figure 3: Attention Distribution of MARVEL-DPR. The attention weights of image features are shown in Figure 3(a). And the attention weight entropy of image captions and features is shown in Figure 3(b).
Finetuning Strategies. We then show the effectiveness of different finetuning strategies. In this experiment, we individually finetune the language model (T5) and visual module (CLIP) to show the changes of attention distributions and analyze the behaviors of different finetuning strategies.
As shown in Figure 3. The attention scores are calculated by cross attentions from the decoder to the encoder module of T5. We first show the attention weight distribution of image features in Figure 3(a). When we only finetune the language model, the attention heads tend to allocate more attention weights to image features, helping to adapt the visual module in the language model. On the other hand, the image features win more attention weights when the CLIP module is finetuned. However, as shown in Figure 3(b), only finetuning the CLIP module shows a scattered attention weight mechanism than other models, which misleads the T5-ANCE to capture more important information from encoded representations of documents. All these phenomena demonstrate the necessity of the training strategies of MARVEL, which pretrain visual module for adaption and only finetune the language model for multi-modal retrieval. In addition, we show the retrieval effectiveness with different finetuning methods in Appendix A.2.
Model MRR@10 NDCG@10 Rec@100
MARVEL-DPR 55.71 52.94 88.23
w/ 1-NN Token 38.80 35.89 73.59
w/ 5-NN Tokens 42.39 39.27 75.04
w/ Random Token 37.73 35.34 71.92
MARVEL-ANCE 65.15 62.95 92.40
w/ 1-NN Token 51.37 48.27 80.47
w/ 5-NN Tokens 52.22 49.35 81.88
w/ Random Token 44.22 41.55 71.23
Table 5: Multi-Modal Retrieval Performance of Different Image Feature Replacement Strategies. We conduct experiments on MARVAL-DPR and MARVAL-ANCE models by replacing the image features with the average of $k$-NN ($k$ Nearest Neighbour) word embeddings. The $k$ is set to 1 and 5.
5.4 Learned Semantics of Image Features
In this experiment, we explore the semantic information of image features encoded by the visual module on the WebQA dataset. During training MARVEL model, we map the encoded image features into the input space of T5-ANCE’s word embeddings. We conduct several experiments by replacing the encoded image features with the embeddings of the nearest or random tokens.
As shown in Table 5, replacing encoded image features with $k$-NN token embeddings generally outperforms the retrieval model using randomly selected token embeddings. It demonstrates that the visual plugin module effectively maps image semantics in the input space of the language model, and the ability to keep growing with more token embeddings (5-NN). However, the retrieval performance significantly drops when employing $k$-NN token embeddings to replace the image features, compared to the MARVEL model. It demonstrates the role of encoded image features beyond the semantic representations of word embeddings. The encoded image features may act as a kind of prompt, encouraging language models to capture image semantics Merullo et al. (2023). More cases are shown in Appendix A.4.
6 Conclusion
This paper proposes Multi-modAl Retrieval via Visual modulE pLugin (MARVEL). MARVEL integrates a visual plugin module to a well-trained dense retriever and pretrains the visual module with image-caption contrastive training for adaption. Our MARVEL model achieves state-of-the-art on all benchmarks by unifying the multi-modal document encoding and alleviating the modality gap between images and texts.
Limitations
Even though MARVEL shows strong effectiveness in the multi-modal retrieval task, there are some limitations in our work. Existing multi-modal retrieval systems still highly depend on the semantics of image caption instead of the image understanding ability of the visual module. In this case, MARVEL pretrains the visual understanding module but achieves limited improvements. Building an effective visual understanding module is crucial for the multi-modal retrieval task.
Appendix A Appendix
A.1 Experimental Details of MARVEL Pretraining Data
In this subsection, we introduce the experimental details to process the pretraining data.
To pretrain the visual module in MARVEL, we collect the image-caption pairs from the ClueWeb22 dataset. We retain the English pages, extract the content within the image tag and use the image and alt-text to construct the image-caption pair. To ensure the quality of the pretraining dataset, following LAION-400M Schuhmann et al. (2021), we use CLIP to calculate the embeddings of images and captions and compute the cosine similarity between the two embeddings. Subsequently, we discard all samples with a cosine similarity lower than 0.3. The pretraining dataset contains 1.6M image-caption pairs, and we randomly select 10,000 pieces of data as the development set and use the rest for pretraining visual module.
Finetune Modaltiy MRR@10 NDCG@10 Rec@100
Text 53.99 48.36 85.70
CLIP & T5 Image 60.91 62.22 93.71
Multi 55.41 53.02 88.48
Text 54.28 48.24 85.80
T5 Image 60.81 61.74 93.55
Multi 55.58 52.86 88.50
Text 49.39 42.04 74.90
CLIP Image 55.18 55.16 87.91
Multi 44.54 41.03 75.91
Text 48.38 41.39 74.57
N/A Image 51.60 51.63 83.49
Multi 44.09 40.65 74.16
Table 6: The Retrieval Performance with Different Training Strategies. We freeze each module of MARVEL-DPR to explore the benefits of training between different modules.
A.2 Retrieval Effectiveness of Different Finetuning Strategies
In this experiment, we show the performance of single/cross and multi-modal retrieval tasks with different finetuning strategies.
As shown in Table 6, finetuning the CLIP module indeed improves the retrieval performance of the whole frozen model, especially in the image retrieval task. This observation shows that multi-modal training signals are effective to benefit the capability of visual modules. When we only tune the parameters of T5, MARVEL-DPR achieves significant improvements over the frozen model, showing the language model’s strong ability to adapt the visual module to the dense retriever. Nevertheless, the fully finetuned model decreases the retrieval performance of MARVEL-DPR that only finetunes T5. It shows the necessity of the pretraining-and-then-finetuning strategy of MARVEL, which pretrains the visual understanding module for adaption and finetunes the language model for multi-modal retrieval.
A.3 More Details of ClueWeb22-MM
Data Type Median Average Max Min
Queries 8.0 9.9 245.0 1.0
Text Documents 52.0 127.8 1121183.0 1.0
Image Captions 6.0 8.1 998.0 1.0
Table 7: Length Statistics of Queries, Text Documents and Image Captions in ClueWeb22-MM Dataset.
Range of Image Sizes Number
Height or Width $\geq$ 1024 23.8k
Height and Width $\geq$ 1024 7.4k
Height or Width $\geq$ 512 81.9k
Height and Width $\geq$ 512 43.9k
Height or Width $\geq$ 256 234.6k
Height and Width $\geq$ 256 170.2k
Table 8: Image Size Distribution of ClueWeb22-MM.
Queries with the Text Document as Label
Query: Chinese Dragons — Facts, Culture, Origins, and Art
Text Document: Live updates on China travel restrictions for 2022. Learn more Home Chinese Culture Traditional Chinese Clothes Chinese Dragons — Facts, Culture, Origins, and Art Written by Mike Ho Updated Dec. 14, 2021 Chinese dragons are powerful and benevolent symbols in Chinese culture, with supposed control over watery phenomenon, e.g. summoning rain during a drought. Dragons are everywhere in China — in legends, festivals, astrology, art, names, and idioms.
Query: How to manage partitions with the Disk Management tool, in Windows | Digital Citizen
Text Document: Disk Management A new window should pop up, listing the drive letter of the partition. Click or tap Change and, in the next window, select the new drive letter you wish to assign to it. Then, click or tap OK.
Query: here’s a small-batch peanut butter oatmeal cookie recipe for you
Text Document: You are here: Home / Recipes / Small-batch Peanut Butter Oatmeal Cookies Small-batch Peanut Butter Oatmeal Cookies 02/21/19 | Cookies , Desserts , Recipes , Small-batch Dessert These Small-batch Peanut Butter Oatmeal Cookies are the perfect cookie hybrid. They’re rich and peanut buttery, bendy and chewy, and the best of both worlds. A few weeks ago, I posted these (AMAZING) Peanut Butter Oatmeal Cookies . It was a big-batch recipe meant for sharing and freezing, so I promised that I’d add a small-batch version ASAP for those of you who are here for small-batch desserts. So here we go. Let’s make a cute little batch of Peanut Butter Oatmeal Cookies and share with no one.
Query: What foods increase uric acid
Text Document: Vegetables and legumes that increase uric acid Legumes such as lentils, chickpeas or beans are known for their purine content, so their intake should be limited to only once or twice a week if you have high uric acid. Other vegetables that should be eaten in moderation are asparagus, mushrooms, cauliflower, spinach, radishes and leeks… Other foods that increase uric acid Other foods that increases uric acid and should be avoided are: All kinds of alcoholic beverages , especially beer and wine. Carbonated beverages, sugar-laden soft drinks and packaged juices. Avoid cooking with brewer’s yeast…
Queries with the Image Document as Label
[Uncaptioned image] Query: Use Web apps With the New Chromium Edge on Windows 10 Image Caption: Web Apps Running Chromium Edge
[Uncaptioned image] Query: What are Runestones In Witcher 3? Image Caption: Witcher 3 best runewords
[Uncaptioned image] Query: Everything We Know About Mindy Kaling and BJ Novak’s Relationship—Including Sweet Details from Her Book Image Caption: mindy-kaling-bj-novak-removebg
[Uncaptioned image] Query: Vector Cross Product Formula Excel Template Image Caption: Vector Cross Product Formula-1.2
Table 9: Examples of ClueWeb22-MM. We give practical examples of queries, image documents, and text documents.
Figure Text
[Uncaptioned image] Manual Caption: Mardi Gras Bourbon Street 2015 Bourbon Street, New Orleans, during Mardi Gras
Nearest Tokens: [“kehr”,“voted”, “brightness”, “event”, “city”, “venue”, “local”, “pub”, “bounce”, “island”, “ferry”,
“keto”, “Ice”, “residents”, “lighting”, “store”, “lights”, “banks”, “Lake”, “impacted”, “lively”, “drinks”, “eye”]
Nearest Tokens w/o CLIP Pretraining: [“OUG”, “7,000”, “ban”, “CU”, “edited”, “ition”, “Pop”, “imprisonment”,
“O”, “militari”, “immuno”, “Ton”, “reset”, “États”, “concise”, “Arbeits”, “IN”, “Hi”, “RAM”,
“Hello”, “stocked”, “charged”, “institu”]
[Uncaptioned image] Manual Caption: Red-shanked Douc at the Philadelphia Zoo
Nearest Tokens: [“Whale”, “endangered”, “horn”, “bird”, “goat”, “animals”, “wildlife”, “mammals”, “whale”, “Gib”,
“Elephant”, “Savannah”, “dach”, “birds”, “creatures”, “Wildlife”, “lois”, “biomass”, “limb”, “Creatures”]
Nearest Tokens w/o CLIP Pretraining: [“RAM”, “bilingual”, “MOD”, “native”, “recognizable”, “Graphic”, “charged”,
“ordentlich”, “gray”, “suffisamment”, “colorful”, “clar”, “haunt”, “riad”, “CM”, “ammunition”, “ordre”, “thetic”,
“Hi”, “auftrag”, “he”, “ban”, “sets”, “7,000”, “representation”]
[Uncaptioned image] Manual Caption: Military parade in Baku on an Army Day28 Military parade in Baku on an Army Day
Nearest tokens: [“vehicles”, “Fahrzeug”, “territories”, “flag”, “chemical”, “replies”, “migrants”, “parliament”, “bikes”,
“militari”, “equipment”, “République”, “troops”, “clothing”, “gear”, “prisoners”, “machinery”, “tribe”, “vorgesehen”]
Nearest Tokens w/o CLIP Pretraining: [ “împreună”, “RAM”, “ened”, “troupe”, “Compet”, “sie”, “own”, “RGB”,
“Ha”,“operation”, “arbeit”, “enforcement”, “Cor”, “EU”, “LCD”, “countries”, “SO”, “institu”, “grief”, “limbi”,
“default”, “16”, “raum”, “haunt”, “unanimous”]
[Uncaptioned image] Manual Caption:Parlament Wien Austria, Vienna, Austrian Parliament Building
Nearest Tokens: [“Schloss”, “funds”, “Statut”, “furniture”, “Albany”, “structure”, “palace”, “Capitol”, “statute”, “locul”,
“headquarters”, “occupie”, “structures”, “legislature”, “cinéma”, “legislation”, “governmental”, “Argentin”,“vederea”]
Nearest Tokens w/o CLIP Pretraining: [ “militari”, “reset”, “Ton”, “shrine”, “commands”, “hi”, “împreună”, “Achtung”,
“genug”, “shake”, “RAM”, “iconic”, “committed”, “département”, “colo”, “Hi”, “Sammlung”, “pop”, “1951”, “ban”,
“russia”, “Color”, “vivid”, “HM”, “arbeit”, “default”]
[Uncaptioned image] Manual Caption: Green Corn Dance at Tesuque Indian Pueblo near Santa Fe, New Mexico File name: 06 10 015670
Nearest Tokens: [“parametri”, “damaged”, “members”, “Feuer”, “muddy”, “potentiel”, “membres”, “statutory”, “aires”,
“monetary”, “pun”, “Chemical”, “bookstore”, “Donna”, “tari”, “elemente”, “lawmakers”, “IN”, “bankruptcy”, “tomato”,
“chemical”, “materiale”, “banks”, “souvent”, “graphical”, “paper”, “vorgesehen”]
Nearest Tokens w/o CLIP Pretraining: [ “immediate”,“OR”, “ammunition”, “comune”, “cartoon”, “LAR”, “SO”, “LED”,
“sign”, “Sin”, “unanimous”, “entière”, “auftrag”, “historischen”, “concise”, “Arbeits”, “obligaţi”, “gray”, “vivid”, “Hi”,
“enforcement”, “part”, “UD”, “He”, “suffisamment”, “ened”, “agencies”, “EU”, “charged”, “commune”, “clown”]
Table 10: The Nearest Tokens of Image Features. We randomly select five image documents, encode these image features using the visual module of MARVEL and MARVEL w/o CLIP Pretraining, and then show the nearest tokens of the encoded image features. The tokens related to the semantics of the image document are highlighted.
Model Modality WebQA ClueWeb22-MM
MRR@10 NDCG@10 Rec@100 MRR@10 NDCG@10 Rec@100
MARVEL-ANCE Text 64.72 58.88 90.26 71.73 75.40 92.29
Image 66.12 67.49 95.12 77.57 81.34 96.50
Multi 65.15 62.95 92.4 55.19 62.83 93.16
w/o Image Caption Text 64.67 58.30 89.49 69.75 73.32 90.60
Image 3.85 4.32 24.81 18.26 20.65 45.07
Multi 33.70 30.83 56.45 37.29 40.74 64.26
w/o Image Feature Text 63.42 57.95 90.27 71.17 74.78 91.57
Image 64.32 65.42 94.15 76.83 80.60 95.88
Multi 63.60 61.43 91.99 54.98 62.64 92.60
Table 11: Additional Ablation Study Results on MARVEL-ANCE.
To show the details of our ClueWeb22-MM dataset, we show the data collection, data processing and data statistics.
Data Collection. Following previous work in text retrieval Zhang et al. (2020); Xie et al. (2023), we regard the anchor text as a query and assume that its linked web page is related to the query. Then we extract image documents and text documents from these anchor-linked web pages. To obtain image documents, we parse HTML to extract the content within the image tag, then use alt-text as image caption, and crawl the image features from the image URL.
Data Processing. Ensuring the quality and meaningfulness of the ClueWeb22-MM dataset, we conduct additional processing on the data to filter out noise data according to the quality of images and alt-texts. Concerning images, we retain data with image file extensions such as jpg/png/jpeg and discard samples with image URLs containing keywords, e.g. “logo”, “button”, “icon”, “plugin”, or “widget”. Besides, we exclude the example, which has empty alt-text, has “no alt attribute” and contains an alt-text that is shorter than 5.
To further guarantee the quality of the dataset, we use T5-ANCE Yu et al. (2023) to estimate the relevance between the anchor and its corresponding image document. We encode all captions of image documents using T5-ANCE, use the anchor texts as queries to retrieve the images and reserve the anchors that are ranked in the top 10. Finally, we respectively sample 10,000 queries to build the development set and test set. The rest data are used for finetuning models, which contain 72,028 queries.
Data Statistics. We provide length statistics on queries, text documents, and image captions in Table 7 and present the image size distribution in Table 8. Subsequently, as shown in Table 9, we show eight examples to illuminate the ClueWeb22-MM dataset. These examples show that the anchor-document pairs are of high quality. Thus we can use them to build an evaluation benchmark for multi-modal retrieval.
A.4 Effectiveness of Image Features
We show some case studies in Table 10 to show the effectiveness of the visual understanding module by verbalizing the semantics of encoded image features using word tokens.
We randomly select five image documents of different topics and represent the encoded image feature with some tokens to verbalize the image semantics. Specifically, we first use the visual plugin modules of MARVEL and MARVEL w/o CLIP Pretraining to encode the image features. Then, to show the semantics of the encoded image features, we utilize cosine similarity to find the $k$-NN tokens for each encoded image feature. Finally, we utilize the token with the highest score to represent the semantics of the encoded image features.
Overall, compared with the MARVEL w/o CLIP Pretraining model, MARVEL learns more effective representations for image documents. These representations are closer to the word tokens associated with the semantics of the image documents. It demonstrates that MARVEL has the ability to better adapt the visual understanding module to dense retrievers.
For the first two examples, MARVEL w/o CLIP Pretraing learns more similar representations for both image documents. The related word tokens of these image documents contain lots of same tokens, such as “7,000”, “Hi”, and “RAM”, which are unrelated to the semantics of the image documents. On the contrary, our MARVEL model can learn more similar semantics to both image documents. Specifically, MARVEL verbalizes the first image document using the words “brightness”, “resident” and “store”, which are related to the image description of “Bourbon Street”. And MARVEL also learns the semantics of “animals”, “wildlife” and “creatures” of the second image document. The next two instances show the effectiveness of MARVEL in learning more fine-grained semantics of the image documents by verbalizing the image documents with more related words, such as “militari”, “vehicle”, “flag”, “legislatur”, and “government”. All these cases illustrate our visual module pretraining strategy’s important role in improving MARVEL’s visual understanding ability.
Even though our visual module pretraining method shows its effectiveness, the last case shows the limitation of existing CLIP based models. The image document describes the semantics of “Green Corn Dance”, which is still hard to comprehend by the visual understanding module.
A.5 Additional Ablation Studies on MARVEL
We conduct additional ablation studies to explore the role of image captions and image features in the multi-modal retrieval task.
As shown in Table 11, the relevance modeling between queries and image documents heavily depends on the image caption, which is also observed in previous work Liu et al. (2023b). The image features contribute to approximately 1% improvements in the image retrieval task, demonstrating the effectiveness of image features in helping model better understand the image documents.
(a) Top5 Multi-modal Documents Retrieved from WebQA.
(b) Top5 Multi-modal Documents Retrieved from Clueweb22-MM.
Figure 4: Case Studies. We present two cases from WebQA and ClueWeb22-MM and show the top5 retrieved multi-modal documents. The ground-truth documents and related content are highlighted in red and blue respectively.
A.6 Case Studies
In Figure 4, we show two cases from WebQA and ClueWeb22-MM to study the multi-modal retrieval effectiveness of MARVEL. The top 5 documents retrieved by UniVL-DR, MARVEL-DPR, and MARVEL-ANCE are presented.
For the first case, UniVL-DR conducts shallow keyword matching and returns text documents that are related to “animal” and “Peace” mentioned in the query, which are unrelated to the query. MARVEL can better understand that “Peace and Plenty” is a famous painting and retrieve more related images and text documents for users. In the second case, UniVL-DR, MARVEL-DPR, and MARVEL-ANCE all return images or text documents related to “promotion ideas”. Notable, MARVEL can better understand the user’s question and return the modal that the user expects. MARVEL-ANCE introduces a variety of sales promotion strategies rather than matching on “promotion” keywords. It shows the effectiveness of MARVEL in better fusing the retrieval results from different modalities, which thrives on universal multi-model document encoding.

### v4

@@
 Specifically, we enable the image understanding ability of the well-trained dense retriever, T5-ANCE, by incorporating the visual module’s encoded image features as its inputs.
-To facilitate the multi-modal retrieval tasks, we build the ClueWeb22-MM dataset based on the ClueWeb22 dataset, which regards anchor texts as queries, and exacts the related text and image documents from anchor-linked web pages.
+To facilitate the multi-modal retrieval tasks, we build the ClueWeb22-MM dataset based on the ClueWeb22 dataset, which regards anchor texts as queries, and extracts the related text and image documents from anchor-linked web pages.
 Our experiments show that MARVEL significantly outperforms the state-of-the-art methods on the multi-modal retrieval dataset WebQA and ClueWeb22-MM.
-MARVEL provides an opportunity to broaden the advantages of text retrieval to the multi-model scenario.
+MARVEL provides an opportunity to broaden the advantages of text retrieval to the multi-modal scenario.
 Besides, we also illustrate that the language model has the ability to extract image semantics and partly map the image features to the input word embedding space.
@@
 MARVEL: Unlocking the Multi-Modal Capability of Dense Retrieval via Visual Module Plugin
-[…]
+Tianshuo Zhou1thanks: indicates equal contribution., Sen Mei [^1] footnotemark: 1, Xinze Li1, Zhenghao Liu1thanks: indicates corresponding author., Chenyan Xiong2, Zhiyuan Liu3, Yu Gu1, Ge Yu1 1Department of Computer Science and Technology, Northeastern University, China 2Language Technologies Institute, Carnegie Mellon University, United States 3Department of Computer Science and Technology, Institute for AI, Tsinghua University, China Beijing National Research Center for Information Science and Technology, China
 1 Introduction
@@
 UniVL-DR Liu et al.
-(2023b) encodes queries and multi-modal documents into a universal embedding space for multi-model retrieval.
+(2023b) encodes queries and multi-modal documents into a universal embedding space for multi-modal retrieval.
 However, this work encodes image features and texts using different encoders from CLIP Radford et al.
@@
 The document collection $\mathcal{D}$ not only contains texts $\mathcal{T}=\{d_{\text{Text}}^{1},...,d_{\text{Text}}^{m}\}$, but also includes images =$\mathcal{I}=\{d_{\text{Image}}^{1},...,d_{\text{Image}}^{n}\}$.
-The multi-model retrieval task requires retrievers to conduct relevance modeling, cross-modal matching, and modality fusion Liu et al.
+The multi-modal retrieval task requires retrievers to conduct relevance modeling, cross-modal matching, and modality fusion Liu et al.
 (2023b).
@@
 Different from previous work Liu et al.
-(2023b), we can universally encode query $q$ and multi-model documents $\mathcal{D}=\{d_{\text{Text}}^{1},...,d_{\text{Text}}^{m},d_{\text{Image}}^{1},...,d_{\text{Image}}^{n}\}$ using one encoder, T5-ANCE-CLIP:
+(2023b), we can universally encode query $q$ and multi-modal documents $\mathcal{D}=\{d_{\text{Text}}^{1},...,d_{\text{Text}}^{m},d_{\text{Image}}^{1},...,d_{\text{Image}}^{n}\}$ using one encoder, T5-ANCE-CLIP:
 $\begin{split}\vec{q}&=\text{T5-ANCE-CLIP}(q);\\ \vec{d}_{\text{Text}}^{i}&=\text{T5-ANCE-CLIP}(d_{\text{Text}}^{i});\\ \vec{d}_{\text{Image}}^{i}&=\text{T5-ANCE-CLIP}(d_{\text{Image}}^{i}(I),d_{\text{Image}}^{i}(C)),\end{split}$ (2)
@@
 (2021), and conduct the T5-ANCE-CLIP model.
-We can use a universal encoder, T5-ANCE-CLIP, to encode texts, image features and image documents.
+We can use a universal encoder, T5-ANCE-CLIP, to encode texts, image features, and image documents.
 Specifically, we encode the image feature $I$ using the visual encoder of CLIP Radford et al. (2021) and get its encoded visual representation $\vec{h}^{I}$:
@@
 5 Evaluation Result
-In this section, we first evaluate the performance of MARVEL and then conduct ablation studies.
-Subsequently, we explore the effectiveness of different visual and language model fusion methods and analyze the role of image features in the MARVEL.
+In this section, we first evaluate the performance of MARVEL and conduct ablation studies.
+Then, we explore the effectiveness of different visual and language model fusion methods and analyze the role of visual module adaption pretraining in MARVEL.
 Some case studies are shown in Appendix A.8.
 5.1 Overall Performance
-The multi-modal retrieval performance of MARVEL and baseline models is shown in Table 2.
+The multi-modal retrieval performance of MARVEL and baseline models is shown in Table 2. Besides retrieval performance, we also compared retrieval efficiency in Appendix A.7.
 Overall, MARVEL significantly outperforms baseline models on all datasets by achieving more than 2% improvements on both datasets, demonstrating its advantages in multi-modal retrieval tasks.
@@
 As shown in Table 3, we conduct ablation studies to explore the role of different modules of MARVEL in multi-modal retrieval. More ablation studies are shown in Appendix A.5.
-In the comparison between MARVEL and MARVEL (w/o CLIP Pretraining), pretraining the visual understanding module shows its effectiveness by improving the performance on single/multi-modal retrieval tasks.
+In the comparison between MARVEL and MARVEL w/o CLIP Pretraining, pretraining the visual understanding module shows its effectiveness by improving the performance on single/multi-modal retrieval tasks.
 It shows that the image-caption alignment relations provide some opportunities to adapt the visual module to the language model via pretraining.
-Subsequently, MARVEL also outperforms MARVEL (w/o MS MARCO Pretraining), especially on the text retrieval task.
+Subsequently, MARVEL also outperforms MARVEL w/o MS MARCO Pretraining, especially on the text retrieval task.
 It demonstrates that MARVEL can broaden the advantage of text relevance modeling to the multi-modal retrieval task.
@@
 Text 51.84 85.06
-MARVEL-Sum Image 58.09 93.13
+T5-CLIP (Sum) Image 58.09 93.13
 Multi 35.03 79.00
 Text 48.71 81.78
-MARVEL-Concat Image 37.20 81.14
+T5-CLIP (Concat) Image 37.20 81.14
 Multi 25.19 62.77
 Text 54.28 85.80
-MARVEL Image 60.81 93.55
+T5-CLIP (Plugin) Image 60.81 93.55
 Multi 55.58 88.50
 Table 4: Retrieval Performance of the Models using Different Visual-Language Fusion Methods.
-The Sum and Concat methods are similar to previous work Liu et al.
+T5-CLIP (Sum/Concat) is similar to previous work Liu et al.
 (2023b), which only replace the image caption encoder with T5-ANCE.
@@
 These models have different architectures and are pretrained on text retrieval and image-caption matching tasks.
-The multi-modal retrieval performance of CLIP-Sum decreases when we encode the image caption with a stronger retrieval model T5-ANCE (MARVEL-Sum) instead of CLIP.
+The multi-modal retrieval performance of CLIP-Sum decreases when we encode the image caption with a stronger retrieval model T5-ANCE (T5-CLIP-Sum) instead of CLIP.
 It demonstrates that incorporating an additional visual module into a well-trained dense retrieval model is still challenging for multi-modal retrieval.
@@
 (b) Attention Entropy.
-Figure 3: Attention Distribution of MARVEL-DPR.
+Figure 3: Attention Distribution of MARVEL-ANCE.
 The attention weights of image features are shown in Figure 3(a).
@@
 We first show the attention weight distribution of image features in Figure 3(a).
-When we only finetune the language model, the attention heads tend to allocate more attention weights to image features, helping to adapt the visual module in the language model.
+When we only finetune the language model, the attention heads tend to allocate more balanced attention weights between image features and captions, helping to adapt the visual module in the language model.
 On the other hand, the image features win more attention weights when the CLIP module is finetuned.
@@
 In addition, we show the retrieval effectiveness with different finetuning methods in Appendix A.3.
-Model MRR@10 NDCG@10 Rec@100
-MARVEL-DPR 55.71 52.94 88.23
-w/ 1-NN Token 38.80 35.89 73.59
-w/ 5-NN Tokens 42.39 39.27 75.04
-w/ Random Token 37.73 35.34 71.92
-MARVEL-ANCE 65.15 62.95 92.40
-w/ 1-NN Token 51.37 48.27 80.47
-w/ 5-NN Tokens 52.22 49.35 81.88
-w/ Random Token 44.22 41.55 71.23
-Table 5: Multi-Modal Retrieval Performance of Different Image Feature Replacement Strategies.
-We conduct experiments on MARVAL-DPR and MARVAL-ANCE models by replacing the image features with the average of $k$-NN ($k$ Nearest Neighbour) word embeddings.
-The $k$ is set to 1 and 5.
-5.4 Learned Semantics of Image Features
-In this experiment, we explore the semantic information of image features encoded by the visual module on the WebQA dataset.
-During training MARVEL model, we map the encoded image features into the input space of T5-ANCE’s word embeddings.
-We conduct several experiments by replacing the encoded image features with the embeddings of the nearest or random tokens.
-As shown in Table 5, replacing encoded image features with $k$-NN token embeddings generally outperforms the retrieval model using randomly selected token embeddings.
-It demonstrates that the visual plugin module effectively maps image semantics in the input space of the language model, and the ability to keep growing with more token embeddings (5-NN).
-However, the retrieval performance significantly drops when employing $k$-NN token embeddings to replace the image features, compared to the MARVEL model.
-It demonstrates the role of encoded image features beyond the semantic representations of word embeddings.
-The encoded image features may act as a kind of prompt, encouraging language models to capture image semantics Merullo et al.
-(2023).
-More cases are shown in Appendix A.4.
+Figure Text
+[Uncaptioned image] Manual Caption: Mardi Gras Bourbon Street 2015 Bourbon Street, New Orleans, during Mardi Gras
+Nearest Tokens: [“kehr”,“voted”, “brightness”, “event”, “city”, “venue”, “local”, “pub”, “bounce”, “island”, “ferry”,
+“keto”, “Ice”, “residents”, “lighting”, “store”, “lights”, “banks”, “Lake”, “impacted”, “lively”, “drinks”, “eye”]
+Nearest Tokens w/o CLIP Pretraining: [“OUG”, “7,000”, “ban”, “CU”, “edited”, “ition”, “Pop”, “imprisonment”,
+“O”, “militari”, “immuno”, “Ton”, “reset”, “États”, “concise”, “Arbeits”, “IN”, “Hi”, “RAM”,
+“Hello”, “stocked”, “charged”, “institu”]
+[Uncaptioned image] Manual Caption: Red-shanked Douc at the Philadelphia Zoo
+Nearest Tokens: [“Whale”, “endangered”, “horn”, “bird”, “goat”, “animals”, “wildlife”, “mammals”, “whale”, “Gib”,
+“Elephant”, “Savannah”, “dach”, “birds”, “creatures”, “Wildlife”, “lois”, “biomass”, “limb”, “Creatures”]
+Nearest Tokens w/o CLIP Pretraining: [“RAM”, “bilingual”, “MOD”, “native”, “recognizable”, “Graphic”, “charged”,
+“ordentlich”, “gray”, “suffisamment”, “colorful”, “clar”, “haunt”, “riad”, “CM”, “ammunition”, “ordre”, “thetic”,
+“Hi”, “auftrag”, “he”, “ban”, “sets”, “7,000”, “representation”]
+[Uncaptioned image] Manual Caption: Military parade in Baku on an Army Day28 Military parade in Baku on an Army Day
+Nearest tokens: [“vehicles”, “Fahrzeug”, “territories”, “flag”, “chemical”, “replies”, “migrants”, “parliament”, “bikes”,
+“militari”, “equipment”, “République”, “troops”, “clothing”, “gear”, “prisoners”, “machinery”, “tribe”, “vorgesehen”]
+Nearest Tokens w/o CLIP Pretraining: [ “împreună”, “RAM”, “ened”, “troupe”, “Compet”, “sie”, “own”, “RGB”,
+“Ha”,“operation”, “arbeit”, “enforcement”, “Cor”, “EU”, “LCD”, “countries”, “SO”, “institu”, “grief”, “limbi”,
+“default”, “16”, “raum”, “haunt”, “unanimous”]
+[Uncaptioned image] Manual Caption:Parlament Wien Austria, Vienna, Austrian Parliament Building
+Nearest Tokens: [“Schloss”, “funds”, “Statut”, “furniture”, “Albany”, “structure”, “palace”, “Capitol”, “statute”, “locul”,
+“headquarters”, “occupie”, “structures”, “legislature”, “cinéma”, “legislation”, “governmental”, “Argentin”,“vederea”]
+Nearest Tokens w/o CLIP Pretraining: [ “militari”, “reset”, “Ton”, “shrine”, “commands”, “hi”, “împreună”, “Achtung”,
+“genug”, “shake”, “RAM”, “iconic”, “committed”, “département”, “colo”, “Hi”, “Sammlung”, “pop”, “1951”, “ban”,
+“russia”, “Color”, “vivid”, “HM”, “arbeit”, “default”]
+Table 5: The Nearest Tokens of Image Features.
+We randomly select five image documents, encode these image features using the visual module of MARVEL and MARVEL w/o CLIP Pretraining, and then show the nearest tokens of the encoded image features.
+The tokens related to the semantics of the image document are highlighted.
+5.4 Effectiveness of Visual Module Adaption Pretraining
+We show some cases in Table 5 to show the effectiveness of the visual understanding module by verbalizing encoded image features.
+Besides, more experiments about the encoded image features are shown in Appendix A.6.
+We randomly select four image documents of different topics and represent the encoded image feature with some tokens to verbalize the image semantics.
+Specifically, we first use the visual plugin modules of MARVEL and MARVEL w/o CLIP Pretraining to encode the image features.
+Then, to show the semantics of the encoded image features, we utilize cosine similarity to find the $k$-NN tokens for each encoded image feature.
+Finally, the tokens with the highest score are used to represent the semantics of the encoded image features.
+For the first two examples, MARVEL w/o CLIP Pretraing learns more similar representations for both image documents.
+The related word tokens of these image documents contain lots of same tokens, such as “7,000”, “Hi”, and “RAM”, which are unrelated to the semantics of the image documents.
+On the contrary, our MARVEL model can learn more similar semantics to both image documents.
+Specifically, MARVEL verbalizes the first image document using the words “brightness”, “resident” and “store”, which are related to the image description of “Bourbon Street”.
+Additionally, MARVEL captures the semantics of “animals”, “wildlife” and “creatures” in the second image document.
+The next two instances show the effectiveness of MARVEL in learning more fine-grained semantics of the image documents by verbalizing the image documents with more related words, such as “militari”, “vehicle”, “flag”, “legislatur”, and “government”.
+Overall, compared with the MARVEL w/o CLIP Pretraining model, MARVEL has the ability to learn more effective representations that are closer to the semantics of the images, demonstrating the important role of MARVEL’s visual module pretraining strategy in adapting the visual understanding module for dense retrievals.
 6 Conclusion
 This paper proposes Multi-modAl Retrieval via Visual modulE pLugin (MARVEL).
-MARVEL integrates a visual plugin module to a well-trained dense retriever and pretrains the visual module with image-caption contrastive training for adaption.
+MARVEL integrates a visual plugin module with a well-trained dense retriever and pretrains the visual module with image-caption contrastive training for adaption.
 Our MARVEL model achieves state-of-the-art on all benchmarks by unifying the multi-modal document encoding and alleviating the modality gap between images and texts.
@@
 Building an effective visual understanding module is crucial for the multi-modal retrieval task.
+Acknowledgments
+Tianshuo Zhou, Sen Mei, Xinze Li and Zhenghao Liu are supported by the Natural Science Foundation of China under Grant (No.
+62206042 and No.
+62272093), the Joint Funds of Natural Science Foundation of Liaoning Province (No.
+2023-MSBA-081), and the Fundamental Research Funds for the Central Universities under Grant (No.
+N2416012).
+Yu Gu is supported by the National Natural Science Foundation of China (No.
+U23B2019).
+Ge Yu is supported by the Natural Science Foundation of China under Grant (No.
+62137001).
 Appendix A Appendix
-A.1 Experimental Details of MARVEL Pretraining Data
+A.1 License
+We show the licenses of the datasets that we use.
+WebQA uses CC0-1.0 license, while ClueWeb22 shows its terms of use at website [^3] https://lemurproject.org/clueweb22/.
+All these licenses and agreements permit the use of their data for academic purposes.
+A.2 Experimental Details of MARVEL Pretraining Data
 In this subsection, we introduce the experimental details to process the pretraining data.
@@
 Subsequently, we discard all samples with a cosine similarity lower than 0.3.
-The pretraining dataset contains 1.6M image-caption pairs, and we randomly select 10,000 pieces of data as the development set and use the rest for pretraining visual module.
-Finetune Modaltiy MRR@10 NDCG@10 Rec@100
-Text 53.99 48.36 85.70
-CLIP & T5 Image 60.91 62.22 93.71
-Multi 55.41 53.02 88.48
-Text 54.28 48.24 85.80
-T5 Image 60.81 61.74 93.55
-Multi 55.58 52.86 88.50
-Text 49.39 42.04 74.90
-CLIP Image 55.18 55.16 87.91
-Multi 44.54 41.03 75.91
+The pretraining dataset contains 1.6M image-caption pairs, and we randomly select 10,000 pieces of data as the development set and use the rest for the pretraining visual module.
+Finetune Modality MRR@10 NDCG@10 Rec@100
+Text 64.89 58.71 89.98
+CLIP & T5 Image 64.36 65.41 94.19
+Multi 64.37 61.77 91.78
+Text 64.72 58.88 90.26
+T5 Image 66.12 67.49 95.12
+Multi 65.15 62.95 92.40
+Text 48.38 41.63 75.11
+CLIP Image 56.28 56.17 87.67
+Multi 49.22 45.80 80.42
 Text 48.38 41.39 74.57
-N/A Image 51.60 51.63 83.49
-Multi 44.09 40.65 74.16
-Table 6: The Retrieval Performance with Different Training Strategies. We freeze each module of MARVEL-DPR to explore the benefits of training between different modules.
-A.2 Retrieval Effectiveness of Different Finetuning Strategies
+N/A Image 55.09 54.99 87.26
+Multi 48.12 44.75 79.69
+Table 6: The Retrieval Performance with Different Training Strategies. We freeze each module of MARVEL-ANCE to explore the benefits of training between different modules.
+A.3 Retrieval Effectiveness of Different Finetuning Strategies
 In this experiment, we show the performance of single/cross and multi-modal retrieval tasks with different finetuning strategies.
@@
 This observation shows that multi-modal training signals are effective to benefit the capability of visual modules.
-When we only tune the parameters of T5, MARVEL-DPR achieves significant improvements over the frozen model, showing the language model’s strong ability to adapt the visual module to the dense retriever.
-Nevertheless, the fully finetuned model decreases the retrieval performance of MARVEL-DPR that only finetunes T5.
+When we only tune the parameters of T5, MARVEL-ANCE achieves significant improvements over the frozen model, showing the language model’s strong ability to adapt the visual module to the dense retriever.
+Nevertheless, the fully finetuned model decreases the retrieval performance of MARVEL-ANCE that only finetunes T5.
 It shows the necessity of the pretraining-and-then-finetuning strategy of MARVEL, which pretrains the visual understanding module for adaption and finetunes the language model for multi-modal retrieval.
-A.3 More Details of ClueWeb22-MM
+A.4 More Details of ClueWeb22-MM
 Data Type Median Average Max Min
@@
 Table 9: Examples of ClueWeb22-MM. We give practical examples of queries, image documents, and text documents.
-Figure Text
-[Uncaptioned image] Manual Caption: Mardi Gras Bourbon Street 2015 Bourbon Street, New Orleans, during Mardi Gras
-Nearest Tokens: [“kehr”,“voted”, “brightness”, “event”, “city”, “venue”, “local”, “pub”, “bounce”, “island”, “ferry”,
-“keto”, “Ice”, “residents”, “lighting”, “store”, “lights”, “banks”, “Lake”, “impacted”, “lively”, “drinks”, “eye”]
-Nearest Tokens w/o CLIP Pretraining: [“OUG”, “7,000”, “ban”, “CU”, “edited”, “ition”, “Pop”, “imprisonment”,
-“O”, “militari”, “immuno”, “Ton”, “reset”, “États”, “concise”, “Arbeits”, “IN”, “Hi”, “RAM”,
-“Hello”, “stocked”, “charged”, “institu”]
-[Uncaptioned image] Manual Caption: Red-shanked Douc at the Philadelphia Zoo
-Nearest Tokens: [“Whale”, “endangered”, “horn”, “bird”, “goat”, “animals”, “wildlife”, “mammals”, “whale”, “Gib”,
-“Elephant”, “Savannah”, “dach”, “birds”, “creatures”, “Wildlife”, “lois”, “biomass”, “limb”, “Creatures”]
-Nearest Tokens w/o CLIP Pretraining: [“RAM”, “bilingual”, “MOD”, “native”, “recognizable”, “Graphic”, “charged”,
-“ordentlich”, “gray”, “suffisamment”, “colorful”, “clar”, “haunt”, “riad”, “CM”, “ammunition”, “ordre”, “thetic”,
-“Hi”, “auftrag”, “he”, “ban”, “sets”, “7,000”, “representation”]
-[Uncaptioned image] Manual Caption: Military parade in Baku on an Army Day28 Military parade in Baku on an Army Day
-Nearest tokens: [“vehicles”, “Fahrzeug”, “territories”, “flag”, “chemical”, “replies”, “migrants”, “parliament”, “bikes”,
-“militari”, “equipment”, “République”, “troops”, “clothing”, “gear”, “prisoners”, “machinery”, “tribe”, “vorgesehen”]
-Nearest Tokens w/o CLIP Pretraining: [ “împreună”, “RAM”, “ened”, “troupe”, “Compet”, “sie”, “own”, “RGB”,
-“Ha”,“operation”, “arbeit”, “enforcement”, “Cor”, “EU”, “LCD”, “countries”, “SO”, “institu”, “grief”, “limbi”,
-“default”, “16”, “raum”, “haunt”, “unanimous”]
-[Uncaptioned image] Manual Caption:Parlament Wien Austria, Vienna, Austrian Parliament Building
-Nearest Tokens: [“Schloss”, “funds”, “Statut”, “furniture”, “Albany”, “structure”, “palace”, “Capitol”, “statute”, “locul”,
-“headquarters”, “occupie”, “structures”, “legislature”, “cinéma”, “legislation”, “governmental”, “Argentin”,“vederea”]
-Nearest Tokens w/o CLIP Pretraining: [ “militari”, “reset”, “Ton”, “shrine”, “commands”, “hi”, “împreună”, “Achtung”,
-“genug”, “shake”, “RAM”, “iconic”, “committed”, “département”, “colo”, “Hi”, “Sammlung”, “pop”, “1951”, “ban”,
-“russia”, “Color”, “vivid”, “HM”, “arbeit”, “default”]
-[Uncaptioned image] Manual Caption: Green Corn Dance at Tesuque Indian Pueblo near Santa Fe, New Mexico File name: 06 10 015670
-Nearest Tokens: [“parametri”, “damaged”, “members”, “Feuer”, “muddy”, “potentiel”, “membres”, “statutory”, “aires”,
-“monetary”, “pun”, “Chemical”, “bookstore”, “Donna”, “tari”, “elemente”, “lawmakers”, “IN”, “bankruptcy”, “tomato”,
-“chemical”, “materiale”, “banks”, “souvent”, “graphical”, “paper”, “vorgesehen”]
-Nearest Tokens w/o CLIP Pretraining: [ “immediate”,“OR”, “ammunition”, “comune”, “cartoon”, “LAR”, “SO”, “LED”,
-“sign”, “Sin”, “unanimous”, “entière”, “auftrag”, “historischen”, “concise”, “Arbeits”, “obligaţi”, “gray”, “vivid”, “Hi”,
-“enforcement”, “part”, “UD”, “He”, “suffisamment”, “ened”, “agencies”, “EU”, “charged”, “commune”, “clown”]
-Table 10: The Nearest Tokens of Image Features.
-We randomly select five image documents, encode these image features using the visual module of MARVEL and MARVEL w/o CLIP Pretraining, and then show the nearest tokens of the encoded image features.
-The tokens related to the semantics of the image document are highlighted.
 Model Modality WebQA ClueWeb22-MM
@@
 Image 66.12 67.49 95.12 77.57 81.34 96.50
-Multi 65.15 62.95 92.4 55.19 62.83 93.16
+Multi 65.15 62.95 92.40 55.19 62.83 93.16
 w/o Image Caption Text 64.67 58.30 89.49 69.75 73.32 90.60
@@
 Table 10: Additional Ablation Study Results on MARVEL-ANCE.
-To show the details of our ClueWeb22-MM dataset, we show the data collection, data processing and data statistics.
+Model MRR@10 NDCG@10 Rec@100
+MARVEL-DPR 55.71 52.94 88.23
+w/ 1-NN Token 38.80 35.89 73.59
+w/ 5-NN Tokens 42.39 39.27 75.04
+w/ Random Token 37.73 35.34 71.92
+MARVEL-ANCE 65.15 62.95 92.40
+w/ 1-NN Token 51.37 48.27 80.47
+w/ 5-NN Tokens 52.22 49.35 81.88
+w/ Random Token 44.22 41.55 71.23
+Table 11: Multi-Modal Retrieval Performance of Different Image Feature Replacement Strategies.
+We conduct experiments on MARVAL-DPR and MARVAL-ANCE models by replacing the image features with the average of $k$-NN ($k$ Nearest Neighbour) word embeddings.
+The $k$ is set to 1 and 5.
+Setting Model Encoding Time (ms) Retrieval Time (ms)
+Query Img Doc Text Doc Query
+Single Modality (Text Only) BM25 - 0.1 0.1 41.9
+DPR (Zero-Shot) 1.8 2.7 4.7 48.4
+CLIP-Text (Zero-Shot) 0.4 0.2 0.3 32.7
+Anchor-DR (Zero-Shot) 2.2 2.0 2.0 47.7
+T5-ANCE (Zero-Shot) 2.1 2.0 2.0 48.3
+BERT-DPR 1.1 2.7 3.1 47.6
+NQ-DPR 1.3 2.8 4.9 47.8
+NQ-ANCE 1.3 2.7 4.7 48.1
+Divide-Conquer VinVL-DPR 1.4 5.2 4.7 48.4
+CLIP-DPR 0.4 0.8 0.3 32.8
+BM25 & CLIP-DPR 0.4 0.8 0.1 37.3
+UnivSearch CLIP 0.4 0.8 0.3 32.8
+VinVL-DPR 1.3 5.3 4.7 47.6
+UniVL-DR 0.4 0.9 0.3 32.9
+MARVEL 2.1 3.7 2.0 48.0
+Table 12: Retrieval Efficiency.
+We compare the encoding and retrieval times of different architectural models on the same device.
+These models encode the multi-model documents and queries offline and construct the FlatL2 index using FAISS Johnson et al.
+(2019) for online retrieval.
+To show the details of our ClueWeb22-MM dataset, we show the data collection, data processing, and data statistics in this subsection.
 Data Collection.
@@
 We encode all captions of image documents using T5-ANCE, use the anchor texts as queries to retrieve the images and reserve the anchors that are ranked in the top 10.
-Finally, we respectively sample 10,000 queries to build the development set and test set.
+Finally, we respectively sample 10,000 queries to build the development and test set.
 The rest data are used for finetuning models, which contain 72,028 queries.
@@
 Thus we can use them to build an evaluation benchmark for multi-modal retrieval.
-A.4 Effectiveness of Image Features
-We show some case studies in Table 10 to show the effectiveness of the visual understanding module by verbalizing the semantics of encoded image features using word tokens.
-We randomly select five image documents of different topics and represent the encoded image feature with some tokens to verbalize the image semantics.
-Specifically, we first use the visual plugin modules of MARVEL and MARVEL w/o CLIP Pretraining to encode the image features.
-Then, to show the semantics of the encoded image features, we utilize cosine similarity to find the $k$-NN tokens for each encoded image feature.
-Finally, we utilize the token with the highest score to represent the semantics of the encoded image features.
-Overall, compared with the MARVEL w/o CLIP Pretraining model, MARVEL learns more effective representations for image documents.
-These representations are closer to the word tokens associated with the semantics of the image documents.
-It demonstrates that MARVEL has the ability to better adapt the visual understanding module to dense retrievers.
-For the first two examples, MARVEL w/o CLIP Pretraing learns more similar representations for both image documents.
-The related word tokens of these image documents contain lots of same tokens, such as “7,000”, “Hi”, and “RAM”, which are unrelated to the semantics of the image documents.
-On the contrary, our MARVEL model can learn more similar semantics to both image documents.
-Specifically, MARVEL verbalizes the first image document using the words “brightness”, “resident” and “store”, which are related to the image description of “Bourbon Street”.
-And MARVEL also learns the semantics of “animals”, “wildlife” and “creatures” of the second image document.
-The next two instances show the effectiveness of MARVEL in learning more fine-grained semantics of the image documents by verbalizing the image documents with more related words, such as “militari”, “vehicle”, “flag”, “legislatur”, and “government”.
-All these cases illustrate our visual module pretraining strategy’s important role in improving MARVEL’s visual understanding ability.
-Even though our visual module pretraining method shows its effectiveness, the last case shows the limitation of existing CLIP based models.
-The image document describes the semantics of “Green Corn Dance”, which is still hard to comprehend by the visual understanding module.
 A.5 Additional Ablation Studies on MARVEL
@@
 (2023b).
-The image features contribute to approximately 1% improvements in the image retrieval task, demonstrating the effectiveness of image features in helping model better understand the image documents.
+The image features contribute to approximately 1% improvements in the image retrieval task, demonstrating the effectiveness of image features in helping the model better understand the image documents.
+A.6 Learned Semantics of Image Features
+In this experiment, we explore the semantic information of image features encoded by the visual module on the WebQA dataset.
+During training MARVEL model, we map the encoded image features into the input space of T5-ANCE’s word embeddings.
+We conduct several experiments by replacing the encoded image features with the embeddings of the nearest or random tokens.
+As shown in Table 11, replacing encoded image features with $k$-NN token embeddings generally outperforms the retrieval model using randomly selected token embeddings.
+It demonstrates that the visual plugin module effectively maps image semantics in the input space of the language model, and the ability to keep growing with more token embeddings (5-NN).
+However, the retrieval performance significantly drops when employing $k$-NN token embeddings to replace the image features, compared to the MARVEL model.
+It demonstrates the role of encoded image features beyond the semantic representations of word embeddings.
+The encoded image features may act as a kind of prompt, encouraging language models to capture image semantics Merullo et al.
+(2023).
+A.7 Retrieval Efficiency of MARVEL
+In this section, we compare the retrieval efficiency of MARVEL with other baselines on the same device, as shown in Table 12.
+MARVEL follows the general dense retrieval framework Karpukhin et al.
+(2020) for efficient document retrieval.
+It encodes the entire corpus offline and constructs the document index using FAISS Johnson et al.
+(2019) for online searches.
+While the offline encoding time for queries and image/text documents in MARVEL is longer than that in UniVL and other baseline models, this encoding process does not impact retrieval efficiency.
+When comparing the retrieval latency of these models in retrieving the top 100 relevant documents for each query, MARVEL’s retrieval time is comparable to its base model, T5-ANCE, which has a retrieval time of 48.3 ms.
+Furthermore, the retrieval time for document encoding using the CLIP model is less than that of models such as T5 and BERT.
+This demonstrates that the retrieval time is influenced by the dimensionality of the embeddings, and MARVEL’s architecture does not introduce any additional retrieval latency.
 (a) Top5 Multi-modal Documents Retrieved from WebQA.
@@
 The ground-truth documents and related content are highlighted in red and blue respectively.
-A.6 Case Studies
+A.8 Case Studies
 In Figure 4, we show two cases from WebQA and ClueWeb22-MM to study the multi-modal retrieval effectiveness of MARVEL.
@@
 For the first case, UniVL-DR conducts shallow keyword matching and returns text documents that are related to “animal” and “Peace” mentioned in the query, which are unrelated to the query.
-MARVEL can better understand that “Peace and Plenty” is a famous painting and retrieve more related images and text documents for users.
-In the second case, UniVL-DR, MARVEL-DPR, and MARVEL-ANCE all return images or text documents related to “promotion ideas”.
-Notable, MARVEL can better understand the user’s question and return the modal that the user expects.
+In contrast, MARVEL can better understand that “Peace and Plenty” is a famous painting and retrieve more related images and text documents for users.
+In the second case, UniVL-DR, MARVEL-DPR, and MARVEL-ANCE all return documents related to “promotion ideas”.
+Notable, MARVEL can better understand the user’s query and return the expected modality.
 MARVEL-ANCE introduces a variety of sales promotion strategies rather than matching on “promotion” keywords.
-It shows the effectiveness of MARVEL in better fusing the retrieval results from different modalities, which thrives on universal multi-model document encoding.
+It shows the effectiveness of MARVEL in better fusing the retrieval results from different modalities, which thrives on universal multi-modal document encoding.